# Detecting and Mitigating System-Level Anomalies of Vision-Based Controllers

### v1

@@ I Introduction
In this work, we present an approach to detect and mitigate such system-level anomalies for autonomous systems that leverage learning-driven vision-based controllers for decision making. Our work builds upon the study conducted in [19] that leverages a visual simulator within a Hamilton-Jacobi Reachability framework to expose the closed-loop failures of the system under the vision-based controller. However, the proposed framework requires privileged information about the test environment and is computationally not suitable for online applications. Our key idea is to utilize this framework offline to stress-test and automatically mine the system-level failures of the closed-loop system across a variety of environment conditions. Ultimately, this process provides us with a diverse set of input image sequences which when seen by the vision-based controller leads to an overall system failure. These images are then used to train a simple classifier that can serve as an anomaly detector during runtime. We demonstrate that the resultant AD implicitly leverages the revealed failure modes to classify whether a previously unseen input image has the potential to trigger a system failure, without requiring any domain-specific heuristics. Next, we employ the AD to trigger a simple fallback controller online that can trade the system’s performance to assure system safety whenever the anomaly detector determines the system is at risk of entering the failure zone. We illustrate the joint properties of our anomaly detector and fallback controller on an autonomous aircraft taxiing case study, leveraging a vision-based controller. We compare our method against commonly used component-level AD techniques, such as predicting ensemble-based uncertainty, and prediction error-based anomaly detectors to highlight the key advantages of the proposed system-level anomaly detector.
@@ II Problem Setup
Consider a robot whose dynamics are given by, $\dot{\mathbf{x}}=f(\mathbf{x},u)$ where, state $\mathbf{x}\in\mathbb{R}^{n}$ and control $u\in\mathbb{U}$ (a compact set). The robot possesses a sensor $S$, that allows it to perceive visual inputs from its surroundings at any given state $\mathbf{x}$. We can express the mapping between the state to the input as, $I=S(\mathbf{x})$. $I$ could be an RGB image, a pointcloud, etc. Additionally, we have a vision-based controller, $\uppi$, that maps $I$ to the control $u$, and defined as, $u\coloneqq\uppi(I)$. Note that $\uppi$ can be an end-to-end policy or it can consists of several sub-modules some of which may not be data-driven.
@@
where $1$ means that $I$ is anomalous, and 00 means it is not. Note that an ideal $\sigma$ should output $1$ whenever $I\in\mathcal{I}_{unsafe}$ and 00 otherwise. In addition, we seek to find a mitigation system that can preserve system safety even if the robot encounters an anomalous input in $\mathcal{I}_{unsafe}$.
Generating $\sigma$ requires addressing a few challenges: (a) $\mathcal{I}_{unsafe}$ is typically hard to obtain as the test environment is known a priori; (b) even in known environments, obtaining high-dimensional inputs, e.g., RGB images, that lead to system-level failures is a challenging problem; finally, (c) for vision-based controllers, $\mathcal{I}_{unsafe}$ varies with changes in the robot’s surroundings, e.g., $\mathcal{I}_{unsafe}$ for an indoor navigation robot might change as we move from one room to another. Thus, formulation of a general AD pipeline that can recognize features of anomalous input images, without a prior knowledge of the new surrounding is non-trivial.
@@
Hence, the policy $\uppi$ is a composition of the CNN and the P-Controller. Intuitively, the P-controller is designed to steer the aircraft towards the centreline based on the state estimate provided by the CNN.
The image observations are obtained using the X-Plane flight simulator that can render the RGB image, $I$, from a virtual camera ($S$) mounted on the right wing of the aircraft at any state and a given time of day (see Fig. 1(b)-(g) for representative images under different simulation conditions). Note that the CNN here is also trained in simulation using the data collected from X-Plane.
@@
III Background: Hamilton-Jacobi Reachability Analysis
@@
In reachability analysis, we focus on calculating the Backward Reachable Tube (BRT) of the system. The BRT refers to the collection of initial states from which an agent, starting from these states and following policy $\uppi(\mathbf{x})$, can reach the target set $\mathcal{O}$ within the time interval $[t,T]$:
@@
HJ reachability analysis allows us to compute the BRT for general nonlinear systems, even when dealing with control and disturbance inputs affecting the system within arbitrarily shaped target sets. Specifically, to compute the BRT, the target set is first represented as a sub-zero level set of a function $l(\mathbf{x})$, denoted as $\mathcal{O}=\{\mathbf{x}:l(\mathbf{x})\leq 0\}$ [22, 23]. The function $l(\mathbf{x})$ typically represents the signed distance from a state to the target set $\mathcal{O}$. With this formulation, the BRT computation can be reframed as an optimal control problem that involves finding a value function defined as:
@@
with $\text{with }V(\mathbf{x},T)=l(\mathbf{x})\vspace{-0.75em}$
In this equation, $D_{t}$ and $\nabla$ represent the time and spatial gradients of the value function, respectively. The Hamiltonian, denoted as $H\coloneqq{\left\langle\nabla V(\mathbf{x},t),f(\mathbf{x},\uppi(\mathbf{x})\right\rangle}$, embeds the system dynamics in the HJI-VI. Essentially, (6) is a continuous-time counterpart to the Bellman equation in discrete time. Once the value function is determined, the BRT is obtained as the set of states from which entry into the target set is unavoidable. Consequently, the BRT corresponds to the subzero level set of the value function:
@@ IV-C Fallback controller
$\vspace{-0.25em}u_{\text{filtered}}=\begin{cases}\uppi^{*}(I)),&\text{if }\sigma(I)=1\\ \uppi(I),&\text{otherwise,}\end{cases}$ (9)
@@ V Case Study: Autonomous Aircraft Taxiing
Anomaly Detector. We next train a binary classifier on the collected dataset. Our classifier is a pre-trained EfficientNet-B0 Model, replacing the last layer with a fully connected layer that feeds into a softmax output. We trained the classifier using cross-entropy loss and Adam optimizer for 20 epochs.
@@
Figure 3: Some of the failures detected by AD. (a, b) Images correspond to the aircraft being close to the runway boundaries (highlighted with the magenta bounding boxes).(c, d) The visual controller confuses the runway markings (highlighted with the cyan bounding boxes) with the centerline and ultimately leads to a system failure. (e, f) Image (f) is (accurately) not classified as an anomaly during the night time (the same image is classified as anomaly during the day, shown in (e)), as the runway lights (highlighted with the yellow bounding boxes) help the visual controller to predict its position accurately and thereby avoid failure.
@@ V-1 Prediction Error-based Labels
We use a prediction error-based labeling scheme instead of the proposed HJ Reachability-based scheme to collect the training dataset for our AD. Specifically, if the TaxiNet prediction error is above a certain threshold for a particular image, then we label this image as an anomaly. We overlay the prediction error-based anomaly labels (blue) on top of the BRT-based labels (red) for one of the environments in Fig. V-1(a).
It is evident that the prediction error-based labels may not be a good representative of the system-level failures. For example, the states near the Yellow star are anomalous as per prediction error but do not actually cause the system failure (the yellow trajectory in Fig. V-1(c)), resulting in a pessimistic AD and hampering the system performance. On the other hand, certain states and images may have a small error from the TaxiNet module perspective (states near the Green star) and are not classified as anomalies; yet, they cascade to a system failure (the green trajectory in Fig. V-1(c)). This results in an overly optimistic nature of the prediction of error-based AD in certain regions. This “unpredictable” nature of the prediction error-based AD persists with the change in the error threshold used for labeling (Fig. V-1(b)). Unsurprisingly, we observe the same phenomenon in the AD trained on these prediction-error-based labels and, thus, omit those results for brevity. This leads us to conclude that component-level, prediction error-based labels may not be a good representative for determining system-level failures.
@@ V-2 Ensembled Predictions
Figure 5: Labels generated using ensembling denoting failures (red) and success (blue).
@@
Fallback Mechanism. Equipped with a capable AD, we designed a simple fallback mechanism to ensure our aircraft’s safety under anomalous inputs. If at any point in its trajectory, the aircraft observes an image $I$ that is classified as an anomaly (i.e., $\sigma(I)=1$), the linear velocity of the plane ($v$ in Eqn. (2)) is reduced by $0.01m/s$. Intuitively, this results in slowing the aircraft every time it encounters an anomaly, ultimately coming to a complete stop if it continues to encounter anomalies.
@@
As soon as the aircraft detects an image that is non-anomalous, the default TaxiNet controller takes over. Fig. 6 shows trajectories of the aircraft under the resultant controller (red trajectories). The nominal trajectories without the fallback controller are shown in dashed black. In the first case (Fig. 6 (a)), the anomaly is triggered due to the aircraft approaching too close to the runway boundary, while in the second case (Fig. 6 (b)) the anomaly is triggered due to semantic failure of the runway markings. In both cases, the TaxiNet controller leads the system off the runway. On the other hand, the fallback controller decreases the aircraft velocity whenever an anomaly is triggered to ensure system safety. This can also be seen from the velocity variation along the red trajectories. To further illustrate the advantage of the fallback controller, we compute the system BRT under the Taxinet controller and the fallback controller pipeline for one of the test cases on KMWH runway (Fig.6(c)). The BRT (i.e., the set of failure states) is much smaller under the safety pipeline (the BRT volume decreases from 28.23% to 16.86%), showing that the proposed mechanism significantly reduces the number of closed-loop system failures.
Failure Modes of AD. Even with all these exciting discoveries, the AD fails to detect some anomalies. We find that this happens particularly for the images corresponding to the states near the BRT boundary. For example, in Fig. 7(b) and 7(c), we show two visually similar images in our test dataset (corresponding to the green and yellow stars in Fig. 7(a)).
@@
Even though the two images are visually similar, one image is inside the BRT and leads to the system failure (hence an anomaly), while the other does not. Such similar images with minor differences are hard to detect for our AD. Finally, we noticed that some of the semantics of the test environment that cause anomalies are not present in the training dataset and are not predicted well by the proposed AD (Fig. 7(d)). Such issues highlight the need for a continual update of anomaly detectors as more data about the system anomalies is obtained.
Incremental Training. A complementary approach to employing an AD can be to perform targeted incremental training of the TaxiNet model on the collected anomalous data. Essentially, these labeled anomalies represent scenarios where TaxiNet fails to perform optimally. Therefore, by conducting specialized training on these failure instances, we aim to fortify TaxiNet’s robustness in handling such cases. We present some preliminary result of incremental training of TaxiNet in Fig. 8. We show the unmodified TaxiNet in grey and the incrementally trained version in blue. The incrementally trained version has significantly fewer failures. This is also evident from the system trajectory from the same starting state under the two controllers.
@@
Even with such promising results, it’s important to acknowledge that incremental training demands a deep understanding of the base network’s training process, such as hyperparameters, dataset augmentation, among others. Moreover, incremental training may suffer from catastrophic forgetting. In contrast, our approach to training a relatively simple AD, as demonstrated in our work, entails significantly less overhead. Nevertheless, it is essential to appreciate the potential of incremental training as a valuable avenue to improve the performance of the vision-based controller which we will explore further in future work.
@@ VI Discussion and Future Work
In this work, we present an approach aimed at identifying and mitigating system-level anomalies of autonomous systems that rely on vision-based controllers for decision-making. By leveraging insights from reachability analysis, our approach learns an anomaly detector that effectively tackles concerns related to system-level safety during runtime. The learned anomaly detector is combined with a fallback controller to significantly reduce the potential for catastrophic system failures.
However, it was seen that the learned AD was unable to recognize failure cases that were near the BRT boundary. To further enhance the performance of the AD, one can take advantage of temporal information, such as the visual history instead of only the current image. Alternatively, instead of relying on a classifier, other approaches could explore unsupervised methods like clustering via support vector machines or k-means applied to our annotated dataset. These techniques aim to capture and analyze common characteristics found in anomalous images and present a promising avenue for future research. Finally, we will further explore leveraging targeted retraining and incremental training on the anomalous dataset to improve the performance of the vision-based controller.

### v3

@@ I Introduction
Figure 1: Detecting and mitigating system-level anomalies of vision based controller. Offline (top row): We compute the Backward Reachable Tube (BRT) of the vision-based system using Hamilton-Jacobi Reachability analysis [19]. This allows us to create a labeled dataset of safe and unsafe images without any manual intervention. We train the anomaly detector using this dataset. Online (bottom row): We run the system on previously unseen environments where the trained anomaly detector triggers a fallback controller or the default controller depending on whether the observed image is anomalous.
In this work, we present an approach to detect and mitigate such system-level anomalies for autonomous systems that leverage learning-driven vision-based controllers for decision making (Fig. 1). Our work builds upon the study conducted in [19] that leverages a visual simulator within a Hamilton-Jacobi Reachability framework to expose the closed-loop failures of the system under the vision-based controller. However, the proposed framework requires privileged information about the test environment and is computationally not suitable for online applications. Our key idea is to utilize this framework offline to stress-test and automatically mine the system-level failures of the closed-loop system across a variety of environment conditions. Ultimately, this process provides us with a diverse set of input image sequences which when seen by the vision-based controller leads to an overall system failure. These images are then used to train a simple classifier that can serve as an anomaly detector during runtime. We demonstrate that the resultant AD implicitly leverages the revealed failure modes to classify whether a previously unseen input image has the potential to trigger a system failure, without requiring any domain-specific heuristics. Next, we employ the AD to trigger a simple fallback controller online that can trade the system’s performance to assure system safety whenever the anomaly detector determines the system is at risk of entering the failure zone. We illustrate the joint properties of our anomaly detector and fallback controller on an autonomous aircraft taxiing case study, leveraging a vision-based controller. We compare our method against commonly used component-level AD techniques, such as predicting ensemble-based uncertainty, and prediction error-based anomaly detectors to highlight the key advantages of the proposed system-level anomaly detector.
@@ II Problem Setup
Consider a robot with dynamics, $\dot{\mathbf{x}}=f(\mathbf{x},u)$ where, state $\mathbf{x}\in\mathbb{R}^{n}$ and control $u\in\mathbb{U}$ (a compact set). The robot possesses a sensor $S$, that allows it to perceive visual inputs from its surroundings at any given state $\mathbf{x}$. We can express the mapping between the state to the input as, $I=S(\mathbf{x})$. $I$ could be an RGB image, a pointcloud, etc. Additionally, we have a vision-based controller, $\uppi$, that maps $I$ to the control $u$, and defined as, $u\coloneqq\uppi(I)$. Note that $\uppi$ can be an end-to-end policy or it can consists of several sub-modules some of which may not be data-driven.
@@
where $1$ means that $I$ is anomalous, and 00 means it is not. Hence, an ideal $\sigma$ should output $1$ whenever $I\in\mathcal{I}_{unsafe}$ and 00 otherwise. In addition, we seek to find a mitigation system that can preserve system safety even if the robot encounters an anomalous input in $\mathcal{I}_{unsafe}$.
Generating $\sigma$ requires addressing a few challenges: (a) $\mathcal{I}_{unsafe}$ is typically hard to obtain as the test environment is unknown a priori; (b) even in known environments, obtaining high-dimensional inputs, e.g., RGB images, that lead to system-level failures is a challenging problem; finally, (c) for vision-based controllers, $\mathcal{I}_{unsafe}$ varies with changes in the robot’s surroundings, e.g., $\mathcal{I}_{unsafe}$ for an indoor navigation robot might change as we move from one room to another. Thus, formulation of a general AD pipeline that can recognize features of anomalous input images without prior knowledge of the new surroundings is non-trivial.
@@
Hence, the policy $\uppi$ is a composition of the CNN and the P-Controller. Intuitively, the P-controller is designed to steer the aircraft towards the centreline based on the state estimate provided by the CNN [20]. We are interested in analyzing this given controller without modifying its underlying design.
The image observations are obtained using the X-Plane flight simulator that can render the RGB image, $I$, from a virtual camera ($S$) mounted on the right wing of the aircraft at any state and a given time of day (see Fig. 2(b)-(g) for representative images under different simulation conditions).
@@
III Background: Hamilton-Jacobi Reachability
@@
In reachability analysis, we focus on calculating the Backward Reachable Tube (BRT) of the system. The BRT refers to the collection of initial states from which an agent following policy $\uppi(\mathbf{x})$, can reach the target set $\mathcal{O}$ within the time interval $[t,T]$:
@@
HJ reachability analysis allows us to compute the BRT for general nonlinear systems. To compute the BRT, the target set is first represented as a sub-zero level set of a function $l(\mathbf{x})$, denoted as $\mathcal{O}=\{\mathbf{x}:l(\mathbf{x})\leq 0\}$ [22, 23]. The function $l(\mathbf{x})$ typically represents the signed distance from a state to the target set $\mathcal{O}$. With this formulation, the BRT computation can be reframed as an optimal control problem that involves finding a value function defined as:
@@
with $\text{with }V(\mathbf{x},T)=l(\mathbf{x})\vspace{-1em}$
In this equation, $D_{t}$ and $\nabla$ represent the time and spatial gradients of the value function, respectively. The Hamiltonian, denoted as $H\coloneqq{\left\langle\nabla V(\mathbf{x},t),f(\mathbf{x},\uppi(\mathbf{x})\right\rangle}$, embeds the system dynamics in the HJI-VI. Consequently, the BRT corresponds to the subzero level set of the value function:
@@ IV-C Fallback controller
$\vspace{-0.25em}u_{\text{filtered}}=\begin{cases}\uppi^{*}(I),&\text{if }\sigma(I)=1\\ \uppi(I),&\text{otherwise,}\end{cases}$ (9)
@@ V Case Study: Autonomous Aircraft Taxiing
Anomaly Detector. We next train a binary classifier on the collected dataset that takes a 224x224x3 input image and returns the probability of the image being anomalous. We use a pre-trained EfficientNet-B0 Model [25] as our backbone, replacing the last layer with a fully connected layer that feeds into a binary softmax output layer. We trained the classifier using cross-entropy loss and Adam optimizer for 20 epochs with a fixed learning rate of $3e^{-4}$ which took around 1 hour on a Nvidia RTX 3090 GPU.
@@
Figure 4: Failures detected by AD. (a, b) Images correspond to the aircraft being close to the runway boundaries (highlighted with the magenta bounding boxes).(c, d) The visual controller confuses the runway markings (highlighted with the cyan bounding boxes) with the centerline and ultimately leads to a system failure. (e, f) Image (f) is (accurately) not classified as an anomaly during the night time (the same image is classified as anomaly during the day, shown in (e)), as the runway lights (highlighted with the yellow bounding boxes) help the visual controller to predict its position accurately and thereby avoid failure.
@@ V-1 Prediction Error-based Labels
We use a prediction error-based labeling scheme instead of the proposed HJ Reachability-based scheme to collect the training dataset for our AD. Specifically, if the TaxiNet prediction error is above a certain threshold for a particular image, then we label this image as an anomaly. We overlay the prediction error-based anomaly labels (blue) on top of the BRT-based labels (red) for one of the environments in Fig. 5(a).
@@
It is evident that the prediction error-based labels may not be a good representative of the system-level failures. For example, the states near the Yellow star are anomalous as per prediction error but do not actually cause the system failure (the yellow trajectory in Fig. 5(c)), resulting in a pessimistic AD and hampering the system performance. On the other hand, certain states and images may have a small error from the TaxiNet module perspective (states near the Green star) and are not classified as anomalies; yet, they cascade to a system failure (the green trajectory in Fig. 5(c)). This results in an overly optimistic nature of the prediction of error-based AD in certain regions. This “unpredictable” nature of the prediction error-based AD persists with the change in the error threshold used for labeling (Fig. 5(b)). Unsurprisingly, we observe the same phenomenon in the AD trained on these prediction-error-based labels and, thus, omit those results for brevity. This leads us to conclude that component-level, prediction error-based labels may not be a good representative for determining system-level failures.
@@ V-2 Ensembled Predictions
Figure 6: Labels generated using ensembling denoting failures (red) and success (white).
@@
Fallback Mechanism. Equipped with a capable AD, we designed a simple fallback mechanism to ensure our aircraft’s safety under anomalous inputs. If at any point in its trajectory, the aircraft observes an image $I$ that is classified as an anomaly (i.e., $\sigma(I)=1$), the linear velocity of the plane ($v$ in Eqn. (2)) is reduced by $0.01m/s$. Intuitively, this results in slowing the aircraft every time it encounters an anomaly, ultimately coming to a complete stop if it continues to encounter anomalies. As soon as the aircraft detects an image that is non-anomalous, the default TaxiNet controller takes over. Fig. 7 shows trajectories of the aircraft under the resultant controller (red trajectories). The nominal trajectories without the fallback controller are shown in dashed black. In the first case (Fig. 7 (a)), the anomaly is triggered due to the aircraft approaching too close to the runway boundary, while in the second case (Fig. 7 (b)) the anomaly is triggered due to semantic failure of the runway markings. In both cases, the TaxiNet controller leads the system off the runway. On the other hand, the fallback controller decreases the aircraft velocity whenever an anomaly is triggered to ensure system safety. This can also be seen from the velocity variation along the red trajectories. To further illustrate the advantage of the fallback controller, we compute the system BRT under the Taxinet controller and the fallback controller pipeline for one of the test cases on KMWH runway (Fig.7(c)). The BRT (i.e., the set of failure states) is much smaller under the safety pipeline (the BRT volume decreases from 28.23% to 16.86%), showing that the proposed mechanism significantly reduces the number of closed-loop system failures.
@@
Finally, with the inclusion of the AD and the fallback controller along with the TaxiNet CNN, the system was able to run at 25Hz as compared to 33Hz when the TaxiNet governs the system as the sole controller. This allows us to run the system online without any severe latency issues.
@@
Failure Modes of AD. Even with all these exciting discoveries, the AD fails to detect some anomalies. We find that this happens particularly for the images corresponding to the states near the BRT boundary. For example, in Fig. 8(b) and 8(c), we show two visually similar images in our test dataset (corresponding to the green and yellow stars in Fig. 8(a)). Even though the two images are visually similar, one image is inside the BRT and leads to the system failure (hence an anomaly), while the other does not. Such similar images with minor differences are hard to detect for our AD. Finally, we noticed that some of the semantics of the test environment that cause anomalies are not present in the training dataset and are not predicted well by the proposed AD (Fig. 8(d)). Such issues highlight the need for a continual update of anomaly detectors as more data about the system anomalies is obtained.
Incremental Training. A complementary approach to employing an AD can be to perform targeted incremental training of the TaxiNet model on the collected anomalous data. Essentially, these labeled anomalies represent scenarios where TaxiNet fails to perform optimally. Therefore, by conducting specialized training on these failure instances, we aim to fortify TaxiNet’s robustness in handling such cases. We present some preliminary result of incremental training of TaxiNet in Fig. 9. We plan on exploring this method further in future work.
@@ VI Discussion and Future Work
In this work, we present an approach aimed at identifying and mitigating system-level anomalies of autonomous systems that rely on vision-based controllers for decision-making. By leveraging insights from reachability analysis, our approach learns an anomaly detector that effectively tackles concerns related to system-level safety during runtime. The learned anomaly detector is combined with a fallback controller to reduce the potential for catastrophic system failures significantly.
In the future, we plan to test our approach over a wide range of real-world scenarios and focus on more intricate fallback mechanisms that can efficiently work for different system-level failures. In addition, to further enhance the performance of the AD, one can take advantage of temporal information, such as the visual history, instead of only the current image. Finally, instead of relying on a classifier, other approaches could explore unsupervised methods like clustering via support vector machines or k-means applied to our annotated dataset.